From owner-hst-status@stsci.edu Wed Apr 1 07:20:31 1998 X-Sender: cwolf@mail.hst.nasa.gov To: george.albright@hq.nasa.gov, sadams@estif2.gsfc.nasa.gov, barnbaum@stsci.edu, preston.m.burch.1@gsfc.nasa.gov, jconwell@ems.jsc.nasa.gov, john.h.campbell.1@gsfc.nasa.gov, bcooke@sssp.hst.nasa.gov, carpenter@tma1.gsfc.nasa.gov, aferro@as.arizona.edu, jeffrey.m.hanley1@jsc.nasa.gov, hst-status@stsci.edu, chenning@pop500.gsfc.nasa.gov, gsfc.hst.general@newsfeed.gsfc.nasa.gov, j.keith.kalinowski.1@gsfc.nasa.gov, ken.ledbetter@hq.nasa.gov, j_lesko@swales.com, jlegg@stx.com, jleibee@hst.nasa.gov, dleckrone@hst.nasa.gov, ann.c.merwarth.1@gsfc.nasa.gov, jmccullo@ems.jsc.nasa.gov, gmorrow@hst.nasa.gov, mgw@resource-intl.com, mnolan@pop100.gsfc.nasa.gov, Jan.Owings@gsfc.nasa.gov, kpulkkinen@v2pop.hst.nasa.gov, gilbert.roth@hq.nasa.gov, rrasche@as.arizona.edu, arankin@pop400.gsfc.nasa.gov, steven.l.smith1@jsc.nasa.gov, stdesk@eso.org, rthompson@as.arizona.edu, edward.weiler@hq.nasa.gov, vweyers@pop400.gsfc.nasa.gov, howard.j.wood.1@gsfc.nasa.gov Subject: Daily Report #2092 04/01/98 HUBBLE SPACE TELESCOPE DAILY REPORT #2092 PERIOD COVERED: 0000Z (UTC) 03/31/98 - 0000Z (UTC) 04/01/98 Daily Status Report as of 091/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7827 (The Structure of Circumbinary Material in Multiple T-Tauri Systems) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class system RW-Aur to study the structure of its circumbinary material. Recent studies of T-Tauri populations in nearby star forming clouds have shown that the majority of stars exist in multiple {binary or higher} systems. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.6 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of CFHT-PL-18, MH06 and HH18. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Four Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC/1/2/3 7448 (Star Formation in a Protogalaxy Candidate at z=2.72) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the star formation region within the protogalaxy candidate MS1512+36 which is located at a redshift of Z = 2.72 . Recently a new population of galaxies at high redshift has been identified by ultraviolet absorption lines which most likely arise from very young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.10 Completed STIS/CCD 7674 (IR Standards) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard source P330E to establish a set of absolute flux standards on the WD scale that extend the existing wavelength coverage to 10300A with the STIS G750L grating setting. The proposal completed with no reported problems. 1.11 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.12 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-1961 and NGC-1672 far from their centers by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.14 Completed STIS/CCD/MA1/MA2 7917 (Effect of MAMA Charge Offsetting on Sensitivity and Dispersion Accuracy) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was operated in order to conserve charge on the STIS MAMA microchannel plates and maintain even detector quantum efficiency. The STIS MAMA spectroscopic modes are differentially offset each month along and across the detector. This 'dithering' of the spectra is a fundamental part of the operation of the MAMAs needed to preserve its lifetime. Thus it is important that we characterize the effect of the offseting on the science data. No problems were encountered. 1.15 Completed Three Sets of WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.16 Completed WF/PC-2/STIS/CCD/NIC/1/2 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One and Two) were used to make observations of the rich star clusters H14-Cen and NGC-2209 which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 17 Successful : 16 The acquisition at 091/0609Z failed during a ZOE. Analysis of the event has not been completed, and no HSTAR has been written as of the time of this report. The effects on science will be detailed in tomorrow's report. Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 22 Successful : 22 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH